on a map 1 inch represents 20 miles. the distance between 2 towns is 6 1/5. how many miles are actually between the two towns?
Answer by checkley77(12844)

1/20=6.2/X CROSS MULTIPLY
X=6.2*20
X=124 MILES BETWEEN THESE 2 TOWNS.